OpenStudy (anonymous):

modulus of 13^2^2011 division 16 plissss plisss :(

OpenStudy (anonymous):

what do you mean by devision 16?

OpenStudy (anonymous):

modulus of 13^2^2011 divided by 16 plissss plisss :(

OpenStudy (anonymous):

its zero if im right

OpenStudy (anonymous):

you mean \[13^{2^{2011}}\] right? If so, then you can simplify it as \[13^{4022}\] because you multiply exponents when raising an exponent to an exponent. now find the modulus of smaller powers of 13 and see if a pattern arises 13mod(16) = 13 13^2mod(16) = 9 13^3mod(16) = 5 13^4mod(16) = 1 13^5mod(16) = 13 13^6mod(16) = 9 So there are four unique remainders for powers of 13 divided by 16 so divide the exponent by four\[4022/4 = 1005 \mod 2\] Thus, 13^{4022} is the equivalent to 13^{2} so its remainder is 9

OpenStudy (anonymous):

owh I'm clear now. ^_^